OpenStudy (anonymous):

What is the solution set of |2x - 3| - 8 = -1? A- 7 or -7 B- 5 or -2 C- 5 or 2 D- No solution

OpenStudy (jdoe0001):

\(\bf |2x - 3| - 8 = -1\implies |2x - 3| = -1+8\\ \quad \\ |2x - 3|=7\implies \begin{cases} +(2x - 3)=7\\ \quad \\ \bf -(2x - 3)=7 \end{cases}\)

OpenStudy (jdoe0001):

so you'd get 2 values for "x"

OpenStudy (anonymous):

Starting from where @jdoe0001 left off, \[2x-3=7\]\[2x=10\]\[x=5\] \[-2x+3=7\]\[-2x=4\]\[x=-2\]

OpenStudy (anonymous):

Thank You Both! :D